Grade 12Physical Chemistry

41Ar is a radioisotope used to measure the flow of gases from smoke chimneys. A sample
of 41Ar decays initially at a rate of 32500 disintegrations/minute. The decay falls to 24500
disintegrations/minute after 75 min. The half-life of the isotope

Punit Sinha
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Dear student,

Rate of disintegration at some time is given as
\dfrac{dN}{dt}=-\lambda N

Hence\dfrac{N}{N_0}=\dfrac{(\dfrac{dN}{dt})_t}{(\dfrac{dN}{dt})_0}=\dfrac{24500}{32500}=\dfrac{49}{65}

Also we know thatN=N_0e^{-\lambda t}
\implies \lambda t=ln(\dfrac{N_0}{N})=ln(\dfrac{65}{49})=\lambda (75)
\implies \lambda =0.00376min^{-1}
Hence half life=t_{1/2}=\dfrac{ln 2}{\lambda}=184min